The ratio of Geoff’s allowance to Ken’s allowance is 7:11. If Ken gives Geoff $84, they will have the same amount of money. How much do they have altogether?

Answer by ikleyn(52780)

.
The ratio of Geoff’s allowance to Ken’s allowance is 7:11.
If Ken gives Geoff $84, they will have the same amount of money.
How much do they have altogether?
~~~~~~~~~~~~~~~~


Ken's allowance 11x  dollars;

Geoff's allowance is 7x dollars.


Had Ken give Geoff $84, the two quantities would be equal


    11x - 84 = 4x + 84.


It is the basic equation for the problem to setup.

Now we should solve it.


    11x - 4x = 84 + 84

       7x    = 168

        x    = 168/7 = 24  dollars.


Altogether, they have  11x + 7x = 11*24 + 7*24 = 432 dollars  and zero cents.    ANSWER

Solved.

--------------

To see many other similar  (and different)  solved problems of this kind,  look into the lesson
    - Virtual exchange between two persons
in this site.